Polarizers An unpolarized beam of intensity lo 413 W/m2 travels in the positive z-direction and is incident from the left on a series of three linear polarizers as shown. The transmission axis of the first polarizer is aligned with the y-axis. The following polarizers make angles of 62 55 degrees, and 63 - 37 degrees with the positive x axis Side Vievw End iew ^y 1) What is 12, the intensity of the beam immediately following the second polarizer? 38.5635798W/m2 Your submissions: 138.5635798 Computed value: 138.5635798 Feedback: Your answer has been judged correct; the exact answer is: 138.563422461699 Submitted: Wednesday, April 18 at 7:35 AM 2) What is lfinal, the intensity of the beam immediately following the third polarizer? 125.3319353 W/m2 Submit Your submissions:125.3319353 Computed value: 125.3319353 Feedback: Your answer has been judged correct; the exact answer is: 125.331814628281 Submitted: Wednesday, April 18 at 7:36 AM

I only need question #5 answered. I just posted the others in case they are helpful in getting the solution to #5.
Answer #1

as unpolarized light passes through the rightmost polarizer,

I1 = I0 / 2

and light becomes polarized along the polarizer axis,


then I = Ii (cos(theta))^2

I2 = I1 (cos(55-37))^2 = (I0/2)(cos18)^2

after that  

I3 = I2 (cos37)^2 = (I0/2) (cos18 cos37)^2


If = I3 = (413/2) (cos18 cos37)^2

If = 119.13 W/m^2 .....Amns